OpenStudy (goformit100):

Let there be any number. let us suppose : 77 How to divide the above accurately in the given ratios like 2:5:8. Help me to know the required theory for it.

OpenStudy (anonymous):

Hmmm. I don't know. Lets see what research says lol.

OpenStudy (goformit100):

How ?^

OpenStudy (gorv):

first sum up the given ratio =2+5+8=15 now \[for 2-------> \frac{ 2 }{ 15 }*77\]

OpenStudy (gorv):

\[\frac{ 5 }{ 15 }*77 and \frac{ 8 }{ 15 } *77\]

OpenStudy (gorv):

@goformit100